Random sample of size 25 from a normally distributed population with the mean of 4 and variance of 16. Find the point estimate of the population mean if P(X>x)=0.9798.
1
Expert's answer
2021-09-12T17:48:12-0400

"n=25 \\

\mu=4 \\

\sigma^2 = 16 \\

\sigma=4 \\

P(X>x) = 0.9798 \\

P(X≤x) = 1 -0.9798 = 0.0202 \\

P(Z≤ \frac{x- 4 }{4 / \sqrt{25}}) = 0.0202 \\

\frac{x- 4 }{4 / 5} = -2.049 \\

x -4= -1.64 \\

x = 4-1.64=2.36"